(Page 1)  Last modified 10/01/04

Subaru Telescope

National Astronomical Observatory of Japan

Semester  \underline{S05B}
Proposal ID \underline{S05B-}
Received  \underline{/ /}

Application Form for Telescope Time

1. Title of Proposal

Cosmological Expansion Measurements using SNIa in elliptical galaxies

2. Principal Investigator

Name:

Doi  Mamoru

Institute: Institute of Astronomy, School of Science, University of Tokyo Mailing Address: 2-21-1, Osawa, Mitaka, Tokyo E-mail Address: doi@ioa.s.u-tokyo.ac.jp

Phone: \underline{0422-34-5084}  Fax:  \underline{0422-34-5041}

3. Scientific Category

Solar System  Normal Stars  Extrasolar Planets  Star and Planet Formation Compact Objects and SNe  Milky Way  Local Group  ISM Nearby Galaxies  Starburst Galaxies  AGN and QSO Activity  QSO Absorption Lines and IGM Clusters of Galaxies  Gravitational Lenses  High-z Galaxies  Deep Surveys Large-Scale Structure  ⋆ Cosmological Parameters  Miscellaneous

4. Abstract (approximately 200 words)

5. Co-Investigators

Name  Institute  Name  Institute

Naoki Yasuda  University of Tokyo Tadayuki Kodama  NAOJ Tomoki Morokuma  University of Tokyo Naohiro Takanashi  University of Tokyo Koichi Tokita  University of Tokyo  Saul Perlmutter  LBNL Chris Lidman  ESO Supernova Cosmology Project Team ***To be added later ***

6. List of Applicants’ Related Publications (last 5 years)

Lidman,C. et al. 2005, A&A, 430, 843, ”Spectroscopic confirmation of high-redshift supernovae with the ESO VLT” Tegmark,M. et al. 2004, PhRvD, 69,10,103501 ”Cosmological parameters from SDSS and WMAP” Kodama,T. et al., 2004, MNRAS, 350, 1005, ”Down-sizing in galaxy formation at z 1 in the Subaru/XMM-Newton Deep Survey (SXDS)” Doi, et al., 2003, IAUC, 8120, ”Supernovae 2002km-2002ky” Knop, R. A. et al., 2003, ApJ, 598, 102, ”New Constraints on Omega_{M}, Omega_{Λ}, and w from an Independent Set of 11 High-Redshift Supernovae Observed with the Hubble Space Telescope” Sullivan, M. et al., 2003, MNRAS, 340, 1057, ”The Hubble diagram of type Ia supernovae as a function of host galaxy morphology” Miyazaki, S. et al., 2002, PASJ, 54, 833 ”Subaru Prime Focus Camera – Suprime-Cam” Mazzali, P. A. et al., 2002, ApJ, 572, 61 ”The Type Ic Hypernova SN 2002ap” Doi,M. et al., 2001, IAUC, 7649, ”Supernovae 2001cq, 2001cr, 2001cs, 2001ct, 2001cu, 2001cv, 2001cw”

Pages 2 and 3 will be used for technical review by observatory staff. Please provide here clear and detailed information for these purposes. The entire proposal including scientific justification will be passed to support astronomers for preparation of observations upon acceptance.

7. Title of Proposal

Cosmological Expansion Measurements using SNIa in elliptical galaxies

8. Observing Run

Instrument  # Nights  Moon  Preferred Dates Acceptable Dates Observing Modes

Suprime-Cam 5  Dark  Oct–Dec  Sep–Jan imaging

Total Requested Number of Nights  5  Minimum Acceptable Number of Nights  3

9. List of Targets (Use an additional sheet if this space is not sufficient)

I do not want observatory staff to see the target names for the technical review.

Target Name  RA  Dec  Equinox  Magnitude (Band)

RXJ0152.7-1357  01 52 42.0  -13 57 52  J2000.0  i = 24.5

RXJ0848.9+4452  08 48 46.9  +44 56 22  J2000.0  i = 24.5

RDCSJ0910+5422  09 10 44.9  +54 22 09  J2000.0  i = 24.5

SXDS fields  02 18 00.0  -05 00 00  J2000.0  i = 24.5

10. Scheduling Requirements

Sequential observations are neccesary to get SN light curves; We prefer 1 night in Oct.(or the 1st month), 2 nights in Nov.(2nd month), and 2 nights in Dec.(third month). If possible, we prefer 10 half nights rather than 5 full nights in order to get well sampled light curves under possible bad weather conditions.

Request Remote Observation

⋆

11. Instrument Requirements 12. Experience

13. Backup Proposal in Poor Conditions (specify object names)

14. Observing Method and Technical Details

Please describe in detail about instrument configuration, exposure time, required sensitivity, and so on.

15. Condition of Closely-Related Past Observations

Please fill in here, if this proposal is a continuation of (or inextricably related with) the previously accepted proposals. This is to describe what kind of relevant/similar proposals have existed in the past and how such previous observations were carried out.

Proposal IDTitle (may be abbreviated)  Observational condition Achievement (%)

16. Post-Observation Status and Publications

Please report the status or outcome of your main Subaru observations carried out in the past. All observations relevant to this proposal (e.g., those enumerated in the above entry 15) must be included here; otherwise, only those within last 3 years suffice.

Year/Month Proposal ID  PI name  Status: data reduction/analysis  Status: publication

17. Thesis Work

This proposal is linked to the thesis preparation of

18. Subaru Open Use Intensive Programs

This is a proposal for Intensive Programs.

Cosmological Expansion Measurements using SNIa in elliptical galaxies. M. Doi et al.

The signature goal of the most ambitious cosmology projects being designed or built this decade (LSST/LST/Panstarrs, Dark Energy Camera, JDEM/SNAP/Destiny, and the South Pole Telescope and other SZ experiments) is the detailed, accurate measurement of the universe’s expansion history, from deceleration through acceleration, to look for clues of the properties and identity of dark energy. Of the small handful of known measurement techniques (SN Ia, cluster counts, S-Z, weak lensing, and baryon oscillations), only Type Ia supernovae (SNe Ia) have actually been developed to the point of routine use. Initial studies of the decelerating universe using SNe at z >_{∼} 1 by both the Higher-Z Team (Riess et al. 2004) and the Supernova Cosmology Project (Fadeyev, Aldering et al. 2004) clearly point to the limiting factor for both statistical and systematic uncertainties: extinction correction of the host galaxy.

One of possible systematic errors come from host galaxy dust extinction, as was discussed in previous studies of nearby galaxies (e.g. Altavilla et al. 2004, Reindl et al. 2005). To be free from possible systematics due to dust extinction or due to luminosity dependence of SNe Ia on host galaxy properties, we propose to make a clean SNIa sample whose host galaxy is early type galaxies with small extinction. How problematic is the extinction correction uncertainty at z >_{∼}1? The correction for the extinction of SNe from dust in the host galaxies is currently the single dominant source of both statistical and systematic error for SNe distances and the derived cosmological parameters – dramatically so at z > 1 even with HST (see Figure 1b). [restate this for Subaru?] The typical color uncertainties for HST-studied z > 1 SNe is 0.08 – 0.1 in B − V, leading to uncertainties in extinction correction (after accounting for intrinsic color uncertainty) of >0.4 mag! This dispersion grows worse, σ ≈ 0.5, after accounting for the uncertainty in the dust reddening coefficient, R_{B} ≡ A_{B}/E(B − V), which Draine (2003) notes can vary from the fiducial value 4.1 by ±0.5. (Note that the actual dispersion about the Hubble-line fit for z > 1 SNe Ia corrected for extinction matches this 0.5 mag value.)

These large dispersions in extinction correction have been dealt with, e.g. in Riess et al. (2004), by applying a strong Bayesian prior to the distribution, assuming knowledge of the dust and SN distribution in the z > 1 host galaxies (shaded contour of Fig. 2b). However, such Bayesian priors are necessarily one-sided (no negative reddening) and hence are known to introduce systematic biases when the error bars are larger at high-redshift than low-redshift (Perlmutter et al. 1999). This bias can be seen in Fig. 2b as the difference between the long-dashed contour and the solid contour. This approach to the extinction analysis is also subject to other obvious sources of systematic biases, for example if the mean value of R_{B} drifts from low to high redshift, as shown by the short-dashed contour of Fig. 2b. How is this problem solved using SNe Ia in ellipticals?

In Sullivan et al. (2003), we showed that the dispersion (including ground-based measurement error) about the Hubble diagram for elliptical-hosted SNe is 0.16 mag — three times smaller than just the measurement uncertainty for extinction-corrected SNe Ia at z > 1 — primarily due to the absence of dust. Thus, SNe Ia in ellipticals are statistically each worth nine times that of SNe in spirals when making cosmological measurements – and without the aforementioned systematics associated with extinction correction. How is it known that dust is not an issue in z >_{∼} 1 cluster ellipticals? Although evidence for dust is found in about 50% of nearby elliptical galaxies, the quantity of dust is generally very small and confined to a central disk where its cross-section is very small. The clearest line of evidence that dust has little effect on stars in elliptical galaxies comes from the tightness of the color-magnitude relation. The dispersion in the colors of early-type galaxies has long been known to be very small in clusters ranging from Coma to intermediate redshifts (Bower et al. 1992; Ellis et al. 1997; Stanford, Eisenhardt & Dickinson 1998; van Dokkum et al. 2001; Blakeslee et al. 2003). Recent results from HST imaging show the same strikingly small dispersion in color extends to redshifts z >_{∼} 1. Why is Suprime-Cam the right instrument to use to generate a sample of SNe at z >_{∼} 1 in elliptical hosts?

There are number of systematic surveys of SNIa for nearby and intermediate redshifts (e.g. SDSS-II, SN Factory, ESSENSE, SNLS) which will give such ”clean” SNIa sample eventually. But in order to measure the cosmological expansion over last 10Gyears, we need deeper observations. Although there are several instruments of ground based telescopes which can go deep enough to find distant SNe, only Suprime-Cam can find enough number of SN candidates because of its large field of view. HSTACS can go deeper (Yasuda et al. 2003) but the field of view is about 1/100, and telescope time is very expensive to get a large enough statistical sample.

Our previous studies show that about 5-10 SNe per Suprime-Cam field (field value), and we can get light curves of these just as we did in 2002 as a part of the SXDS observatory project (Doi et al. 2002, Yasuda et al. 2003, Lidman et al. 2004). As we show in Fig.2 we successfully obtained light curves of more than 15 high-z SNe, and we got a preliminary Hubble diagram (fig.3) for 5 SNe Ia, and also about 10 SNe Ia whose host redshift will be soon observed with Keck and VLT.

(fig.2, 3 light curves and revised Hubble diagram of 2002fall SNe from Keck draft.) In this proposal we propose to carry out rolling SN Ia searches in the same SXDS fields plus several distant clusters of galaxies where we can find many early type galaxies. Our target clusters are all rich clusters at redshift of 0.8-1.2. Since the ratio of early type galaxies is significantly higher than the standard field value, we expect about 2-3 SNe Ia with early-type host per field, and about 10 SNIa in total. We will choose i-band (z-band) for clusters of z∼0.8(z∼1.2) for light curve observations, and we will take z-band (i-band) images for a couple of epochs to obtain color information of SNe.

In order to identify the object and in order to get redshift, we are requesting follow-up telescope time of Keck, VLT, and HST. However, even without this supplementary time, we can use photometric redshifts obtained from Suprime-Cam multi-color images, many of which were already taken as a part of PISCES project (Kodama et al. 2005). We already have images in four bands (VRi’z’) of RXJ0152.7-1357, five bands (BVRi’z’) for RXJ0848.9+4452, and two bands (VR) for MS1054.4-0321. We also will carry out SN searches in SXDS where we have deep photometry in five bands (BVRi’z’) as well as spectroscopic or X-ray data for many galaxies. The images taken here will be also used to study galaxies in clusters and in field environments. References

Altavilla,G., et al., 2004, MNRAS, 349, 1344
Kodama,T. et al, 2005, astro-ph/0502444
Reindl,B., Tammann,G.A., Sandage.A., and Saha,A., 2005, astro-ph/0501664
Sullivan,M., et al., 2003, MNRAS, 340, 1057

Figure 1: (a) Left Panel: The SCP SN Ia Hubble diagram broken into host galaxy types from Sullivan et al. (2003). The SNe in elliptical hosts (filled red circles) show significantly less dispersion, σ = 0.16 mag, including measurement error. (This ground-based measurement error for this z ∼ 0.5 sample is quite close to the HST measurement error at z > 1 in this proposal.) (b) Right Panel: The comparison of the Hubble diagram, before and after extinction correction, for a mixture of SNe Ia in all host types shows the dramatic increase in error bars due to the uncertainty in B − V color being multiplied by R_{B} ≈ 4 and by the uncertainty in R_{B}. The data shown is from the SCP (Knop et al. 2003) and the Riess et al. 2004 GOODS search samples. For the SNe at redshifts z > 1 this yields an uncertainty of ∼0.5 mag, which is consistent with the measured dispersion of 0.5 mag. The ratio of this dispersion to the elliptical-hosted dispersion of panel (a) makes the elliptical-hosted SNe each worth 9 of the extinction-corrected others.

Figure 2: (a) Left Panel: Simulated 68% confidence region on w′ vs w_{0} for the current literature SN sample but with underlying cosmology (w_{0} = −1; w′ = 0). The parameters are poorly constrained because color errors are magnified by R_{B} ≈ 4. (b) Middle Panel: The solid red contour shows reduced uncertainties (excluding systematic bias) using a Baysian prior on the extinction distribution prior to suppress color errors. If the errors are larger at high z than at low z (as with the actual data), this introduces systematic biases. The filled gray contour is from Riess et al. 2004 using this prior. The short-dashed contour shows that this approach is also sensitive to shifts in R_{B} with redshift; the example shifts from 4.1 to 2.6. (c) Right Panel: The goal of this proposal is shown as a confidence region for a simulated new sample of ∼ 10 z >_{∼} 1 SNe Ia found in cluster ellipticals, together with 5 in ellipticals from the past and ongoing GOODS searches, as well as 120 SNe Ia in ellipticals at the lower redshifts now being produced by the ground-based CFHT SN Legacy Survey, the CTIO Essence survey, and (at z < 0.1) the Nearby SN Factory. A SN Hubble diagram in ellipticals avoids the large statistical error problem of panel (a) and the large systematics problem of panel (b).

Figure 3: FOCAS SPECTRUM HERE.

Figure 4: The simulated data set for this proposal, with signal-to-noise at a given redshift and SN epoch based on our previous SN from ACS and NICMOS. The simulated data was fit with our lightcurve analysis program to test the cadence feasibility. We obtain typical errors of 0.07 to 0.13 mag for 0.9 < z < 1.5, including the in the lightcurve timescale stretch correction uncertainty. The bars and symbols at top show the observing time period and scheduled observations for each cluster (with different cadences depending on the cluster z). The same symbols are used for the observations on the lightcurves, to show where a SN might be discovered and followed in its cluster’s time window. Note that the observations are well spread throughout the year (allowing easy HST scheduling, with flexibility since there are other clusters to study if one is difficult to schedule). There are therefore SNe to be observed in our ground-based observing program at almost any time, in addition to the host galaxies that can be observed any time.